OpenStudy (butterflyhope):

Solve: 2x + 19 - 17 + 14x = 2. Please don't give me the answer, I just need help understanding it :) Thanks!

OpenStudy (campbell_st):

start by collecting the like terms 2x + 14x +19 - 17 = 2 now this needs to be simplified. What do you think you would get..?

OpenStudy (butterflyhope):

Well...um 2x and 14x are alike, so I think it would be 2x + 14x, which is 16x. 19 and 17 are alike, so it's 19 - 17 = 2?....

OpenStudy (anonymous):

continue...............

OpenStudy (butterflyhope):

that would leave 16x - 2 = 2? I have a feeling I'm working this out wrong. xD

OpenStudy (campbell_st):

thats correct.... now you can start to solve the equation... next step is to add 2 to both sides of the equation, this will remove the negative 2 16x - 2 = 2 + 2 +2 -------------- what would you get...?

OpenStudy (butterflyhope):

16x = 4.

OpenStudy (anonymous):

16x - 2 = 2 ???@campbell_st it should be 16x+2=2

OpenStudy (butterflyhope):

16x + 2 = 2 -2? -2 16x = 0? o.O

OpenStudy (campbell_st):

oops... sorry about that... ok... then I should have subtracted 2 16x + 2 = 2 - 2 -2 ---------- 16x = 0 now you can solve for x.

OpenStudy (campbell_st):

that's now the correct solution

OpenStudy (anonymous):

seriously x=0...............are u sure question is right?

OpenStudy (butterflyhope):

16x = 0 /0 /0 = x = 0??

OpenStudy (butterflyhope):

That's what my book says....wth

OpenStudy (butterflyhope):

x does = 0, my sis even checked o.o .....I DON'T UNDERSTAND ALGEBRA ; -; Anyway thanks y'all :)

OpenStudy (anonymous):

trust me.....something is fishy...........lol what's the answer ur book says?

OpenStudy (butterflyhope):

The answer key says x = 0 too.... o.o

OpenStudy (anonymous):

so it is right....... O.o o.O